Airlines compute the weight of outbound flights using either standard average weights provided by the Federal Aviation Administration (FAA) or weights obtained from their own sample surveys. The FAA standard average weight for a passengers carry-on items (personal items plus carry-on bags) is 16 pounds. Many airline companies have begun implementing fees for checked bags. Economic theory predicts that passengens will respond to the increase in the price of a checked bag by substituting carry-on bags for checked bags. As a result, the mean weight of a passengers carry-on items is expected to increase after the implementation of the checked-bag fee. Suppose that a particular airlines passengers had a mean weight for their carry-on items of 16 pounds, the FAA standard average weight, before implementation of the checked-bag fee. The airline conducts a hypothesis test to determine whether the current mean weight of its passengers carry-on items is more than 16 pounds. It selects a random sample of 67 passengers and weighs their carry-on items. The sample mean is x17.1 pounds, and the sample standard deviation is s-6.0 pounds. The airline uses a significance level of α-.05 to conduct its hypothesis test. The hypothesis test is an upper tail test. The test statistic follows a distribution. The value of the test statistic is
Answer #1

the hypothesis test is upper tailed test

the test statistic follows standard normal distribution i.e. z test. z = xbar- mu/(s/\sqrt{n})

z = (17.1 -16)/0.733

z = 1.5

hence the value of test statistic is1.5

we have given alpha is 0.05 and z at 0.05 is 1.645

we will reject H0 if z >equal to 1.645

here we will accept null hypothesis because the z value is less than z critical value.

The P-Value is .066807.

using the critical value approach null hypothesis is  accepted.because calculated z is less then z critical.

using the p value approach the null hypothesis is accepted

because the calculated p value is greater than 0.05. therefore we can conclude that the mean weight of airline passenger carry on items has increased after the implementation of checked bag free.